OpenStudy (anonymous):

HELP PLEASE! At a wavelength of 270nm, a 0.040M solution of acetone in water has an absorption of 0.64 in a 1.00cm cell. The absorption of a solution of unknown concentration of acetone in water was 0.48 at the same wavelength and in the same cell. What is the concentration of acetone in the unknown?

OpenStudy (anonymous):

0.030M. is it?

OpenStudy (anonymous):

Can you show me how you got that

OpenStudy (anonymous):

first of all i dont know about chemistry :) i just assumed that more the no of moles of solute in solution,more chances are that photons of light will collide with them and more will be adsorpton. so i took adsorption directly proportional to concentration. so ratio of concentrations must be the ratio of adsorptions. [0.48/0.64]=[x/0.03]. solve for x we get that...

OpenStudy (anonymous):

Thanks!

OpenStudy (jfraser):

the technical explanation is covered by a law known as Beer's Law. It says \[Abs = \epsilon *b *C\] where Abs is the absorbance, e is a constant, b is the pathlength, and C is the molar concentration. The e value is unique to each kind of solution and wavelength. The first set of data gives you the info to find e: \[Abs = \epsilon * b * C\] so \[\epsilon = \frac{Abs}{b*C} = \frac{0.64}{(1.00cm)(0.040M)} = 16cm^{-1}M^{-1}\] the second set of info asks for C using a different absorbance, but the same e value \[C = \frac{Abs}{\epsilon * b} = \frac{0.48}{(16cm^{-1}M^{-1})*(1.00cm)} = 0.030\] Quarkine is correct

OpenStudy (anonymous):

thanks buddy for explaining it the right way..although it does feel good to use shortcuts,yet it never hurts to learn the right way along with it :)